OpenStudy (anonymous):

If you have a rectangle with a width of 5 and length of 13 and make a diagonal line in it, would the length of that line be 18?

Parth (parthkohli):

Pythagorean \(\Rightarrow 5^2 + 13^2 = c^2 \)

OpenStudy (lgbasallote):

 use pythagorean to verify \[18 = \sqrt{5^2 + 13^2}?\] \[18 = \sqrt{25 + 169}?\] \[18 = \sqrt{194}\] \[\sqrt{194} = 13.93\]